OpenStudy (anonymous):

Prove the identity: cos4x=1-8sin^2xcos^2x

OpenStudy (anonymous):

I get that it uses double angle identities , but how you go from there?

OpenStudy (welshfella):

Try Cos(2x + 2x) = cos2xcos2x - sin2xsin2x

OpenStudy (anonymous):

How do you proceed from there?

OpenStudy (welshfella):

Now try using the double angle identities

OpenStudy (zaynab123):

cos4x=cos(2x+2x) cos4x=cos^2 2x-sin^2 2x cos4x=cos2xcos2x-sin2xsin2x cos4x=(cos^2x-sin^2x)(cos^2x-sin^2x)-(2sinxcosx)(2sinxcosx) Using sin^2x+cos^2x=1 cos4x=cos^4x-2cos^2xsin^2x+sin^4x-(4sin^2xcos^2x) cos^4x+sin^4x=1 so cos^4x=1-2cos^2xsin^2x-(4sin^2xcos^2x) finally cos^4x=1-8sin^2xcos^2x